The ratio of AB: BC is 3:2. If AC=15 cm, find the length of AB
Answer by stanbon(75887)

The ratio of AB: BC is 3:2. If AC=15 cm, find the length of AB
-----
AB = 3x
BC = 2x
---
Equation:
3x + 2x = 15
5x = 15
x = 3
----
AB = 3x = 3*3 = 9
=====================
Cheers,
Stan H.
==================